Jump to main content.


Drake Chemical

EPA Weekly Situation Reports for September 1999

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Friday, September 24, 1999

FOR DATES: From 7:00 a.m. Friday, September 17, 1999 to 7:00 a.m. Friday, September 24, 1999.

The prime contractor, OHM, continued to demobilize equipment and facilities from the site. OHM is currently working 4 ten-hour days Monday through Thursday.

Site activities during the time frame of this report included the construction of drainage swales north of the soil stockpile, the placement of geo textile fabric and rock and the lotting of government property. The placement of the geo-textile fabric and stone on the swale along the east side of the TDF pad was completed. The drop inlet at the north east corner of the site was cut off to obtain necessary elevation to match swales. The contractor continues to final grade the area east of the ash buildings and place geo-textile fabric and rock. Other site activities were dust control and the shipment of debris and scrap metal. Related site activities included a weekly progress meeting held on Tuesday, September 21, 1999, with representatives from the USACE, and OHM. Major activities for the next week include the completion of the drainage swales and parking areas east of the TDF pad. The receiving and stockpiling of stone for the storm water drainage system. Excess government property will continue to be sorted into lots in the ash buildings.

PREPARED BY: Chad Thompson for Michael Ogden USACE


Friday, September 17, 1999

FOR DATES: From 7:00 a.m. Friday, September 10, 1999 to 7:00 a.m. Friday, September 17, 1999.

The prime contractor, OHM, continued to demobilize equipment and facilities from the site. OHM is currently working 4 ten-hour days Monday through Thursday. Site activities during the time frame of this report included the construction of drainage swales around the soil stockpile, the removal of the perimeter fence north of Myrtle Street, and the lotting of government property. The placement of the geo-textile fabric and stone on the swale on the east side of the site was completed up to the drop inlet but has not yet been tied in to it. Work began on the swale running along the eastern side of the former fly ash building. The contractor continues to final grade the area east of the ash buildings and place geo-textile fabric and rock. Other site activities were dust control and the shipment of debris and scrap metal.

Related site activities included a weekly progress meeting held on Tuesday, September 14, 1999, with representatives from the USACE, and OHM.

Major activities for the next week include the shipment of rental equipment off site. The receiving and stockpiling of stone for the storm water drainage system, the irrigation of the seeded soil pile and dust control will continue. The contractor will continue re-grading the area east of the bottom ash building and placing geo-textile fabric and stone in these areas. The contractor will continue the construction of the swale along the eastern side of the fly ash building to control storm water. Excess government property will continue to be sorted into lots in the ash buildings.

PREPARED BY: Chad Thompson for Michael Ogden USACE


Friday, September 10, 1999

FOR DATES: From 7:00 a.m. Friday, September 3, 1999 to 7:00 a.m. Friday, September 10, 1999.

The prime contractor, OHM, continued to demobilize equipment and facilities from the site. OHM is currently working 4 ten-hour days Monday through Thursday. Due to the holiday weekend this week OHM worked 4 ten-hour days Tuesday through Friday.

Site activities during the time frame of this report included the construction of drainage swales around the soil stockpile, the removal of the perimeter fence north of Myrtle Street, and the lotting of government property. The liner system from the WWTP was disposed of off site and the stone which was on top of it was back filled on site. The placement of the geo-textile fabric and stone on the swale on the west side of the stockpile was completed and tied into the drop inlet at the south east corner of the bottom ash building. Work continues on the swale running along the eastern side of the site. The contractor continues to final grade the area east of the ash buildings and place geo-textile fabric and rock. Other site activities were dust control and the shipment of debris and scrap metal.

Related site activities included a weekly progress meeting held on Tuesday, September 7, 1999, with representatives from the PADEP, USACE, OHM and the City of Lock Haven in attendance.

Major activities for the next week include the shipment of rental equipment off site. The receiving and stockpiling of stone for the storm water drainage system, the irrigation of the seeded soil pile and dust control will continue. The contractor will continue re-grading the area east of the bottom ash building and placing geo-textile fabric and stone in these areas. The contractor will continue the construction of the swale along the eastern side of the site and east of the bottom ash building to control storm water. Excess government property will continue to be sorted into lots in the ash buildings.

PREPARED BY: Chad Thompson for Michael Ogden USACE


Friday, September 3, 1999

FOR DATES: From 7:00 a.m. Friday, August 27, 1999 to 7:00 a.m. Friday, September 3, 1999.

The prime contractor, OHM, continued to demobilize equipment and facilities from the site. OHM is currently working 4 ten-hour days Monday through Thursday. Due to the holiday weekend next week OHM will work 4 ten-hour days Tuesday through Friday.

Site activities during the time frame of this report included the construction of drainage swales around the soil stockpile, the removal of site trailers, and the watering of the seeded area. The contractor received rock for the drainage swales. The placement of the geo-textile fabric and stone on the swale on the south and west sides continued. The contractor continued to disconnect electrical and communications lines on site. The cutting and capping of water and sewer lines below grade. The transport of excess government property from off site storage areas to the former bottom ash building for separation into lots was begun this week. The contractor began re-grading the area north of Myrtle Street where office trailers and equipment was staged. Other site activities were dust control and the shipment of debris and scrap metal.

Related site activities included a weekly progress meeting held on Tuesday, August 31, 1999, with representatives from the PADEP, USACE, OHM and the City of Lock Haven in attendance.

Major activities for the next week include the shipment of rental equipment off site. The receiving and stockpiling of stone for the storm water drainage system, the irrigation of the seeded soil pile and dust control will continue. The contractor will be re-grading the area east of the bottom ash building and placing geo-textile fabric and stone in these areas. The contractor will continue the construction of the swales around the stockpile to control storm water. Excess government property will continue to be staged in the ash buildings. The stone and liner system which was underneath the WWTP will be removed and the area taken to final grade.

PREPARED BY: Chad Thompson for Michael Ogden USACE

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.